Find the absolute inequality corresponding to x<=-22/3 or x>=11/3.
I have been working on the problem and somehow came up with |x-11/6|>=11/6
I was told that that wasn't exactly correct. I really want to see all the steps so i can learn how to do these problems. I dont have a problem when the numbers are whole numbers, such as x<=12, but im really confused on these fractions. If anyone could show my all the steps and the answer I would greatly appreciate it. Thank you!

Answer by stanbon(75887)

Find the absolute inequality corresponding to x<=-22/3 or x>=11/3.
------------
Try a geometrical approach.
On a number line put -22/3 and +11/3
The midpoint between them is (-22/3+11/3)/2 = -11/3/2 = -11/6
----------------
The distance from -11/6 to 11/3 (or to -22/3) is 33/6
-------------------
You want to express all the x-values that are greater than or equal to
33/6 from -11/6
---------------------
That would be |x-(-11/6)|>=33/6
or
|x+(11/6)|>=11/2
-------------
Cheers,
Stan H.